# SS4Rec: Continuous-Time Sequential Recommendation with State Space Models

### v1

@@ 3.4.2. Discretized by Variable Timesteps
For each user-item interaction sequence, $\mathbf{\Delta T}\in\mathbb{R}^{L}$ denotes its time intervals, and $\mathbf{s}\in\mathbb{R}^{P}$ is the timescale parameter to capture different timescales of each latent channel in the data. The discretization step calculates the product of timescale parameter and time intervals among data, i.e., $\mathbf{\Delta}=\mathbf{\Delta T}\cdot\mathbf{s}\in\mathbb{R}^{L\times P}$.
By following the aforementioned process, the time-aware SSM with time-varying parameters is constructed to capture the system dynamics. In practice, we choose S5 to implement the time-aware SSM due to its proven excellence in time sequence modeling tasks. By integrating system dynamics from a temporal perspective, the time-aware SSM layer offers optimized latent representations. This layer facilitates the relation-aware SSM layer in modeling complex and time-varying relationships between items.
3.5. Relation-Aware SSM
@@
3.5.1. Selection Mechanism
To model the evolving relationship between the data points, Mamba let the SSM parameters be functions of the input. This feature is the “selection mechanism” modifying the original SSM to S6. Specifically, its parameters $\mathbf{B}$, $\mathbf{C}$, and $\mathbf{\Delta}$ are data-dependent and are transformed by different linear projections of given inputs $\mathbf{x}\in\mathbb{R}^{B\times L\times D}$:
(7) $\mathbf{\Delta}=Softplus(\mathbf{Parameter}+\mathbf{s}_{\mathbf{\Delta}}(\mathbf{x}))$ $\in\mathbb{R}^{B\times L\times D},$
@@ 3.5.1. Selection Mechanism
(8) $\mathbf{B}=\mathbf{s}_{\mathbf{B}}(\mathbf{x})=Linear(\mathbf{x})\in\mathbb{R}^{B\times L\times P},$
(9) $\mathbf{C}=\mathbf{s}_{\mathbf{C}}(\mathbf{x})=Linear(\mathbf{x})\in\mathbb{R}^{B\times L\times P}.$
@@
The discretized SSM can be obtained by substituting these parameters into Equation 5 and 6. These parameterizations in the selection mechanism make this SSM layer selectively remember the crucial information and ignore the irrelevant information.
3.6. Interpretation of $\mathbf{\Delta}$
@@
(10) $&\text{Time-Aware SSM: }&\mathbf{\Delta}\leftarrow f_{\theta}(\mathbf{\Delta T})\\ &\text{Relation-Aware SSM: }&\mathbf{\Delta}\leftarrow f_{\theta}(\mathbf{X})\\ \end{aligned}\right.$
Compared to the decoupled two SSMs in this work, another direction worth exploring is the design of a unified SSM whose $\mathbf{\Delta}$ is conditioned on both input data and time, i.e., $\mathbf{\Delta}\leftarrow f_{\theta}(\mathbf{X},\mathbf{\Delta T})$.
3.7. Add & Norm
@@
(11) $\mathbf{X}_{l+1}=LayerNorm(SSM(\mathbf{X}_{l})+\mathbf{X}_{l})$
where $\mathbf{X}\in\mathbb{R}^{B\times L\times D}$, $l$ refers to the $l$-th updated layer, and $LayerNorm(\cdot)$ denotes the layer normalization operation.
3.8. Continuous-Time Prediction
Most current sequential recommenders are next-item predictors, including recommenders that utilize temporal information. This is because RNNs and Transformers are naturally designed as sequence models for next-token prediction, and RNN-based or Transformer-based sequential recommenders can only achieve next-item prediction as well. They neglect the specific time of the next interaction which is useful for exploring the time-dependent preferences of users. Unlike the next-item prediction, our recommender implements continuous-time prediction by the continuous-time parameterization of SSM.
After the stacked SS4Rec blocks, the predictor of our SS4Rec calculates the inner product as a recommendation score according to the next time:
(12) $P(v_{next\_time}=v_{i})=\mathbf{x_{i}}^{T}\mathbf{o}$
where $v_{next\_time}$ refers to the interaction item at time $next\_time$, $\mathbf{o}\in\mathbb{R}^{D}$ denotes the last item embedding in the output embedding of the last SS4Rec block, and $\mathbf{x_{i}}\in\mathbb{R}^{D}$ denotes the candidate item embedding. In our model, the learned representation contains the information of the next interaction time, i.e., the time when an item needs to be recommended for the user. Thereby, SS4Rec can select the most appropriate items at the appropriate time for personalized recommendations.
@@ 3.8. Continuous-Time Prediction
(13) $\mathcal{L}=-\sum_{u\in\mathcal{U}}\sum_{i=1}^{N}{y}_{i}log(P(v_{next\_time}=v_{i}))$
@@ 4. Experiments
4.1. Experiments Setting
4.1.1. Baselines.
A variety of representative sequential recommendation methods are chosen to compare with our method. Reflecting the evolutionary trajectory of sequence model development, these methods have transitioned from RNN-based to attention-based mechanisms, with a recent emphasis on the exploration of SSMs. They are divided into three groups as follows:
(1) RNN-based methods:
  • GRU4Rec (Hidasi et al., 2016): It employs GRU to model user sequences and encode user interests into a final state for session-based recommendation, which treats each user’s sequence of clicks as a session.
  • NARM (Li et al., 2017): A GRU-based method employs an encoder-decoder structure to consider both user’s sequential actions and main purpose in a session.
(2) Attention-based methods:
  • SASRec (Kang and McAuley, 2018): This is a classical sequential recommendation method that leverages the Transformer’s self-attention mechanism to process the input sequence.
  • BERT4Rec (Sun et al., 2019): It employs bidirectional self-attention to model user behavior sequences, allowing each item to incorporate both its left and right contexts in the user’s historical behaviors.
  • TiSASRec (Li et al., 2020): This approach introduces discrete time interval embedding in SASRec, with the ability to capture temporal information.
(3) SSM-based methods:
  • Mamba4Rec (Liu et al., 2024): The first Mamba-based sequential recommender, which mainly substitutes the self-attention mechanism with the Mamba layer, in contrast to SASRec.
  • RecMamba (Yang et al., 2024): This is a simplified model that only uses the mamba layer, which can be regarded as S6-only in the Ablation Study.
4.1.2. Datasets.
Table 1. Datasets Statistics.
Dataset Amazon Sports Amazon Video-Games Movielens-1M KuaiRec(Small) KuaiRec(Big)
@@ 4.1. Experiments Setting
To effectively assess the performance of each method, we select five datasets with different characteristics. These datasets vary greatly in domains, sequence length, time span, and sparsity.
@@
  • Amazon (McAuley et al., 2015): Series of datasets are gathered from the Amazon e-commerce platform, comprising a substantial volume of product reviews extracted from the Amazon website. We extract the categories ”video games” and ”sports” for subsequent experiments.
@@
  • Movielens (Harper and Konstan, 2015): A widely used dataset collected from the Movielens website records user ratings of movies seen. The version (MovieLens-1M) used in our experiments contains 1 million user ratings.
@@
  • KuaiRec (Gao et al., 2022): A real-world dataset is collected from the recommendation logs of the video-sharing mobile app Kuaishou. It contains two interaction matrices, the big and the small matrices, where the small matrix is known as the first dataset to contain a fully observed user-item interaction matrix. In this paper, KuaiRec(Big) refers to the big interaction matrix, while KuaiRec(Small) denotes the small one.
@@
These datasets span significant temporal ranges, with Movielens (1995–2003), Amazon (1996–2018), and KuaiRec (2022) reflecting varied stages in the progression of the Internet. They vary greatly in sequence length and sparsity. In terms of average sequence length, the difference between the longest average sequence length (3,305.69 in KuaiRec(Small)) and the shortest average sequence length (8.54 in Amazon Sports) is more than 200 times. In terms of sparsity, Amazon reaches over 99.99%, while the small matrix of kuairec is less than 1%. Detailed dataset statistics are presented in Table 1.
4.1.3. Implementation Details
@@
To assess the efficacy of our sequential recommendation model, we adhere to the standard evaluation protocol used in this field. Specifically, we employ the leave-one-out strategy for testing within the recbole1.0 code framework (Zhao et al., 2021). Hit Ratio@K(HR@K), Normalized Discounted Cumulative Gain@K(NDCG@K), and Mean Reciprocal Rank@K(MRR@K) are utilized to evaluate the recommendation performance. For all metrics, we set $K=10$. Higher values for these metrics signify better ranking accuracy, which indicates a more effective recommendation model. The specific hyperparameter settings are depicted in Table 2, and all experiments are conducted on a single A800 80GB GPU.
Table 2. Hyper-parameters of SS4Rec.
@@
Dataset Amazon Sports Amazon Video-Games Movielens -1M KuaiRec (Small) KuaiRec (Big)
@@
4.2. Overall Comparison
Table 3. Overall performance comparison. The bold indicates the best performance and the underlined indicates the second best performance.
@@ 4.2. Overall Comparison
The comparison results on the five datasets are shown in Table 3. From the results, we have the following significant observations.
Our proposed SS4Rec model outperforms most baseline models across datasets with varying characteristics. Among these baselines, both attention-based and RNN-based methods show comparable performance; Mamba4Rec performs better than the others, except on the KuaiRec(Small) dataset. This suggests that state space models have a superior ability to model sequences and accurately learn user preferences. Furthermore, our model surpasses Mamba4Rec, highlighting the importance of considering time intervals in sequence recommendations.
In scenarios with varying data densities, our model exhibits different boosting effects. SS4Rec shows a more significant improvement in dense datasets, such as Movielens-1M and KuaiRec(Big), compared to sparser datasets like Amazon Sports and Amazon Video-Games. This is because, in denser scenarios, the model can extract a robust data representation from richer information, enhancing its ability to identify underlying patterns. In KuaiRec(Small), Bert4Rec achieves optimal performance on NDCG@10 and MRR@10, a result attributed to its bi-directional attention mechanism capturing a greater abundance of supervised learning signals in such a highly dense dataset.
In summary, our method, SS4Rec, outperforms all baseline approaches across all five datasets when evaluated using the HR@10 metric. Additionally, it achieves the highest performance on three datasets—Sports, Movielens-1M, and KuaiRec(Big)—based on metrics NDGC@10 and MRR@10. For the KuaiRec(Small) and Video-game datasets, it almost achieves the second-best performance for metrics NDGC@10 and MRR@10. These experimental results demonstrate the effectiveness of SS4Rec in sequential recommendations.
4.3. Time-dependent Prediction
To clearly illustrate SS4Rec’s ability for continuous-time prediction, we construct a Toy dataset wherein user interactions with items are exclusively time-dependent. For instance, a user might typically listen to the morning news at 6 a.m. At this specific time, the recommendation system should be able to identify this time-dependent preference and suggest relevant content to the user. The construction of the Toy dataset is illustrated as follows in Algorithm 1.
@@ 4.3. Time-dependent Prediction
Unlike existing discrete-time recommendation systems that encode time information through discrete time embeddings, SS4Rec is a continuous-time parametric recommendation system. In SS4Rec, time information is encoded using state space functions, allowing the system to capture the dynamic changes of state over time. This continuous-time approach offers higher temporal resolution and flexibility, enabling a more precise understanding of changes in user behavior and preferences. To highlight the superiority of our method among time-aware methods, we introduce temporal information into the comparison baselines. Specifically, we introduce a time interval embedding, which is a $D$-dimensional embedding obtained by a single-layer perceptron based on the value of the input time interval. This embedding is then combined with the item embedding.
@@
The results shown in Figure 2 indicate that SS4Rec outperforms methods that utilize discrete-time embeddings. The use of discrete-time embeddings proves insufficient for capturing the dynamics of user interests. In contrast, SS4Rec effectively models the continuously evolving relationships between input points with irregular time intervals in sequences, thus improving time-dependent recommendations.
[…]
@@
Figure 2. Experiments of Toy Dataset. Time-interval embedding is added to item embedding for baselines.
4.4. Simulation of Partial Observation
[…]
Figure 3. Operation of dropping some items in sequence.
@@ 4.4. Simulation of Partial Observation
In real-world scenarios, users often interact with multiple platforms simultaneously. Consequently, the interaction sequence collected may only be a partial observation of the user’s complete interaction history. To simulate this situation, the items in each sequence are randomly dropped with a 10% probability in training, as shown in Figure 3. We conduct experiments on randomly dropped sequences to assess the performance of sequential recommendation methods in partially observable sequences. The results are shown in Table 4.
@@
Compared to RNN-based and attention-based baselines, SSM-based methods consistently outperform them by a considerable margin on most datasets. This demonstrates the effectiveness of using SSM in sequential recommenders. Our proposed model, SS4Rec, achieves the best or near-best results demonstrating its adaptability to real-world recommendation scenarios. Even with some observations missing, the model is still capable of utilizing updated time intervals to capture the time variation in sequence. For instance, if Item 3 is not observed, the time interval between Item 2 and Item 4 is updated to $\Delta_{2}+\Delta_{3}$. As a result, SS4Rec is better equipped to model the evolution of user preferences over continuous time, enhancing recommendation performance despite partial observations. Additionally, it is noted that Mamba4Rec performs better on short-sequence (such as Sports and Video Games), while SS4Rec excels on sequences with longer average lengths and richer temporal information.
@@
Table 4. Performance of partially observed sequences. We randomly drop 10% items for each sequence.
@@
4.5. Ablation Study
Table 5. Ablation analysis.

### v2

@@ III-D2 Discretization by Variable Timesteps
For each user-item interaction sequence, $\mathbf{\Delta T}\in\mathbb{R}^{L}$ denotes its time intervals between consecutive interactions, and $\mathbf{s}\in\mathbb{R}^{P}$ is the timescale parameter to capture different timescales of each latent channel in the data. The discretization step calculates the product of timescale parameter and time intervals among data, i.e., $\mathbf{\Delta}=\mathbf{\Delta T}\cdot\mathbf{s}\in\mathbb{R}^{L\times P}$.
By following the aforementioned process, the time-aware SSM with time-varying parameters is constructed to capture the system dynamics. In practice, we choose S5 to implement the time-aware SSM due to its proven effectiveness in time sequence modeling tasks. By integrating system dynamics from a temporal perspective, the time-aware SSM layer offers optimized latent representations. This layer facilitates the relation-aware SSM layer in modeling complex and time-varying relationships between items.
III-E Relation-Aware SSM
@@
Selection Mechanism: To model the evolving relationship between the data points, Mamba lets the SSM parameters be functions of the input. This feature is the ”selection mechanism” modifying the original SSM to S6. Specifically, its parameters $\mathbf{B}$, $\mathbf{C}$, and $\mathbf{\Delta}$ are data-dependent and are transformed by different linear projections of given inputs $\mathbf{x}\in\mathbb{R}^{B\times L\times D}$:
$\mathbf{\Delta}=Softplus(\mathbf{Parameter}+\mathbf{s}_{\mathbf{\Delta}}(\mathbf{x}))$ $\in\mathbb{R}^{B\times L\times D},$ (7)
@@
$\mathbf{B}=\mathbf{s}_{\mathbf{B}}(\mathbf{x})=Linear(\mathbf{x})\in\mathbb{R}^{B\times L\times P},$ (8)
$\mathbf{C}=\mathbf{s}_{\mathbf{C}}(\mathbf{x})=Linear(\mathbf{x})\in\mathbb{R}^{B\times L\times P}.$ (9)
@@
The discretized SSM can be obtained by substituting these parameters into (5) and (6). These parameterizations in the selection mechanism make this SSM layer selectively remember the crucial information and ignore the irrelevant information.
III-F Interpretation of $\mathbf{\Delta}$
@@
$&\text{Time-Aware SSM: }&\mathbf{\Delta}\leftarrow f_{\theta}(\mathbf{\Delta T})\\ &\text{Relation-Aware SSM: }&\mathbf{\Delta}\leftarrow f_{\theta}(\mathbf{X})\\ \end{aligned}\right.$ (10)
Compared to the decoupled two SSMs in this work, another direction worth exploring is the design of a unified SSM whose $\mathbf{\Delta}$ is conditioned on both input data and irregular time interval, i.e., $\mathbf{\Delta}\leftarrow f_{\theta}(\mathbf{X},\mathbf{\Delta T})$.
III-G Add & Norm
@@
$\mathbf{X}_{l+1}=LayerNorm(SSM(\mathbf{X}_{l})+\mathbf{X}_{l})$ (11)
where $\mathbf{X}\in\mathbb{R}^{B\times L\times D}$, $l$ refers to the $l$-th updated layer, and $LayerNorm(\cdot)$ denotes the layer normalization funtion.
III-H Continuous-Time Prediction
Most existing sequential recommenders are next-item predictors, including those that utilize temporal information. This constraint arises from the inherent design of RNNs and Transformers as sequence models optimized for next-token prediction. As a result, RNN-based or Transformer-based sequential recommenders are limited to next-item prediction. They neglect the specific timing of the next interaction, which is crucial for capturing users’ time-dependent preferences. Unlike the next-item prediction, our recommender implements continuous-time prediction by the continuous-time parameterization of SSM.
After the stacked SS4Rec blocks, the predictor of our SS4Rec calculates the inner product as a recommendation score based on the next time:
$P(v_{next\_time}=v_{i})=\mathbf{x_{i}}^{T}\mathbf{o}$ (12)
where $v_{next\_time}$ is the interaction item at time $next\_time$, $\mathbf{o}\in\mathbb{R}^{D}$ denotes the last item embedding in the output embedding of the last SS4Rec block, and $\mathbf{x_{i}}\in\mathbb{R}^{D}$ denotes the candidate item embedding. In our model, the learned representation contains the information of the next interaction time, i.e., the time when an item needs to be recommended for the user. Thereby, SS4Rec is capable of selecting the most appropriate items at the optimal time for personalized recommendations.
@@ III-H Continuous-Time Prediction
$\mathcal{L}=-\sum_{u\in\mathcal{U}}\sum_{i=1}^{N}{y}_{i}log(P(v_{next\_time}=v_{i}))$ (13)
@@ IV Experiments
TABLE I: Datasets Statistics.
Dataset Amazon Sports Amazon Video-Games Movielens-1M KuaiRec(Small) KuaiRec(Large)
@@
IV-A Experiments Setting
IV-A1 Baselines
A diverse set of representative sequential recommendation methods is selected for comparison with our method. Reflecting the evolutionary trajectory of sequence model development, these methods have progressed from RNN-based to attention-based mechanisms, with recent advancements exploring SSMs. They are categorized into three groups as follows:
(1) RNN-based methods
  • GRU4Rec [6]: it employs Gated Recurrent Unit (GRU) to model user sequences and encode user interests into a final hidden state for session-based recommendation, where each user’s sequence of clicks is treated as a session.
  • NARM [23]: a GRU-based method adopts an encoder-decoder architecture to capture both user’s sequential actions and primary intent within a session.
(2) Attention-based methods
  • SASRec [10]: this is a classical sequential recommendation method that leverages the Transformer’s self-attention mechanism to process the input sequence.
  • BERT4Rec [11]: it utilizes bidirectional self-attention to model user behavior sequences, allowing each item to incorporated contextual information from both preceding and subsequent interactions in the user’s history.
  • TiSASRec [18]: this approach introduces discrete time interval embedding in SASRec, enabling it to effectively capture temporal information.
(3) SSM-based methods
  • Mamba4Rec [15]: the first Mamba-based sequential recommender, which mainly replaces the self-attention mechanism with the Mamba layer, in contrast to SASRec.
  • RecMamba [16]: this is a simplified model that uses only the mamba layer, which can be referred to as S6-only model in the Ablation Study.
IV-A2 Datasets
@@
To effectively evaluate the performance of each method, we select five datasets with diverse characteristics, differing significantly in terms of domain, sequence length, time span, and sparsity.
@@
  • Amazon [34]: a series of datasets are gathered from the Amazon e-commerce platform, comprising a substantial volume of product reviews extracted from the Amazon website. We extract the categories ”video games” and ”sports” for subsequent experiments.
@@
  • Movielens [35]: a widely used dataset collected from the Movielens website records user ratings of movies seen. The version (MovieLens-1M) used in our experiments contains 1 million user ratings.
@@
  • KuaiRec [36]: a real-world dataset is collected from the recommendation logs of the video-sharing mobile app Kuaishou. It contains two interaction matrices, a large matrix and a small matrix, where the small matrix, referred to as the first dataset, contains a fully observed user-item interaction matrix. In this paper, KuaiRec(Large) refers to the large interaction matrix, while KuaiRec(Small) denotes the small one.
@@
These datasets span significant periods, with Movielens (1995–2003), Amazon (1996–2018), and KuaiRec (2022) reflecting different stages in the evolution of the Internet. They show considerable variation in both sequence length and sparsity. In terms of average sequence length, the difference between the longest average sequence length (3,305.69 in KuaiRec(Small)) and the shortest average sequence length (8.54 in Amazon Sports) is more than 200 times. Regarding sparsity, Amazon exceeds 99.99%, while the small matrix of KuaiRec has a sparsity of less than 1%. Detailed statistics for datasets are provided in Table I.
IV-A3 Implementation Details
@@
To assess the efficacy of our sequential recommendation model, we follow the standard evaluation protocol used in this field. Specifically, we apply the leave-one-out strategy for testing within the recbole1.0 code framework [37]. Hit Ratio@K(HR@K), Normalized Discounted Cumulative Gain@K(NDCG@K), and Mean Reciprocal Rank@K(MRR@K) are utilized to evaluate the recommendation performance, with $K=10$ for all metrics. Higher values for these metrics signify better ranking accuracy, which indicates a more effective recommendation model. The specific hyperparameter settings are depicted in Table II, and all experiments are conducted on a single A800 80GB GPU.
TABLE II: Hyper-parameters of SS4Rec.
@@
Dataset Amazon Sports Amazon Video-Games Movielens -1M KuaiRec (Small) KuaiRec (Large)
@@
IV-B Overall Comparison
TABLE III: Overall performance comparison. The bold indicates the best performance and the underlined indicates the second best performance.
@@ IV-B Overall Comparison
Our proposed SS4Rec model outperforms most baseline models across datasets with diverse characteristics. Among these baselines, both attention-based and RNN-based methods show comparable performance; Mamba4Rec performs better than the others, except on the KuaiRec(Small) dataset. This indicates that state space models are more effective at modeling sequences and learning user preferences. Furthermore, our model outperforms Mamba4Rec, emphasizing the importance of considering time intervals in sequence recommendations.
In scenarios with varying data densities, our model exhibits different levels of improvement. SS4Rec shows more significant gains in dense datasets, such as Movielens-1M and KuaiRec(Large), compared to sparser datasets like Amazon Sports and Amazon Video-Games. This is because, in denser scenarios, the model can extract a robust data representation from richer information, enhancing its ability to identify underlying patterns. In KuaiRec(Small), Bert4Rec achieves optimal performance on NDCG@10 and MRR@10, a result attributed to its bi-directional attention mechanism capturing a greater abundance of supervised learning signals in such a highly dense dataset.
In summary, our method, SS4Rec, outperforms all baseline approaches across all five datasets when evaluated using the HR@10 metric. Additionally, it achieves the highest performance on three datasets—Sports, Movielens-1M, and KuaiRec(Large)—based on metrics NDCG@10 and MRR@10. For the KuaiRec(Small) and Video-Game datasets, it almost achieves the second-best performance for metrics NDCG@10 and MRR@10. These experimental results demonstrate the effectiveness of SS4Rec in sequential recommendations.
IV-C Time-Dependent Prediction
To clearly illustrate SS4Rec’s ability for continuous-time prediction, we construct a Toy dataset wherein user interactions with items are exclusively time-dependent. For instance, a user may typically listen to the morning news at 6 a.m. At this specific time, the recommendation system should be able to identify this time-dependent preference and suggest relevant content to the user. The construction of the Toy dataset is illustrated as follows in Algorithm 1.
@@ IV-C Time-Dependent Prediction
Unlike existing discrete-time recommender systems that encode time information through discrete time embeddings, SS4Rec is a continuous-time parametric recommender system. In SS4Rec, time information is encoded through state space functions, allowing the system to capture the dynamic changes of state over time. This continuous-time approach offers higher temporal resolution and flexibility, enabling a more precise understanding of changes in user behavior and preferences. To highlight the superiority of our method among time-aware methods, we introduce temporal information into the comparison baselines. Specifically, we introduce a time interval embedding, which is a $D$-dimensional embedding obtained by a single-layer perceptron based on the value of the input time interval. This embedding is then combined with the item embedding.
@@
The results shown in Fig. 2 indicate that SS4Rec outperforms methods that utilize discrete-time embeddings. The use of discrete-time embeddings proves insufficient for capturing the dynamics of user interests. In contrast, SS4Rec effectively models the continuously evolving relationships between input points with irregular time intervals in sequences, thus improving time-dependent recommendations.
@@
Figure 2: Experiments on Toy dataset. Time-interval embedding is added to item embedding for baselines.
IV-D Simulation of Partial Observation
Figure 3: Operation of dropping some items in sequence.
@@ IV-D Simulation of Partial Observation
In real-world scenarios, users often interact with multiple platforms simultaneously. Consequently, the interaction sequence collected may only be a partial observation of the user’s complete interaction history. To simulate this situation, the items in each sequence are randomly dropped with a 10% probability in training, as shown in Fig. 3. We conduct experiments on randomly dropped sequences to assess the performance of sequential recommendation methods in partially observable sequences. The results are shown in Table IV.
@@
Compared to RNN-based and attention-based baselines, SSM-based methods consistently outperform them by a considerable margin on most datasets. This demonstrates the effectiveness of using SSM in sequential recommenders. Our proposed model, SS4Rec, achieves the best or near-best results demonstrating its adaptability to real-world recommendation scenarios. Even with some observations missing, the model is still capable of utilizing updated time intervals to capture the time variation in sequence. For instance, if Item 3 is not observed, the time interval between Item 2 and Item 4 is updated to $\Delta_{2}+\Delta_{3}$. As a result, SS4Rec is better equipped to model the evolution of user preferences over continuous time, enhancing recommendation performance despite partial observations. Additionally, it is noted that Mamba4Rec performs better on short-sequence (such as Sports and Video-Games), while SS4Rec excels on sequences with longer average lengths and richer temporal information.
@@
TABLE IV: Performance of partially observed sequences. We randomly drop 10% items for each sequence.
@@
IV-E Ablation Study
We conduct a series of ablation studies to examine each component’s contributions to the proposed architecture. Table V shows the performance of our default method and its five variants.
TABLE V: Ablation analysis.